Operations Report – April 3rd

Crew 279 Operations Report 03-04-2023
SOL: 1
Name of person filing report: Ioana Dimitrova
Non-nominal systems: robotic observatory
Notes on non-nominal systems: awaiting for a solution from the manufacturer
Spirit rover used: no
Beginning charge: NOT MEASURED
Ending charge: NOT MEASURED
Currently charging: YES
Opportunity rover used: no
Beginning charge: NOT MEASURED
Ending charge: NOT MEASURED
Currently charging: YES
Curiosity rover used: yes
Hours: 226.8
Beginning charge: 100%
Ending charge: 92%
Currently charging: yes
Perseverance rover used: yes
Hours: 260.1
Beginning charge: 100%
Ending charge: 91%
Currently charging: yes
General notes on rovers: Spirit and Opportunity not used today and crew forgot to measure them but will be measured tomorrow and the days to come.
Summary of Hab operations: Music influence on stress experiment was conducted. Inventory of food and what everyone ate today for microbiota experiment. Knobs on oven fix (see below for more info).
WATER USE: (see below in questions for mission support) Will be measured tomorrow
Water (static tank): 484.65 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 14:00 to 21:00 for experiments and reports
Summary of suits and radios:
Suits nominal, charging nominal. Radios nominal but technical problem with headset number 5. I think the problem is in the cable of the microphone. Sometimes it works, sometimes not. From other radios, we can always hear that radio number 5 is transmitting something (with the click sound of a walkie talkie), but often you cannot hear the person at all. Which leads us to think the problem is not with the radio itself. We also tried changing headsets with radio number 5, and there is no problem. So the problem must be in the headset, and because sometimes it works and sometimes not, I suppose it’s a cable problem.
We noticed this problem at the beginning of the EVA in the airlock. We aborted the decompression and resolved the problem before decompressing again and going outside. Augustin (who was wearing the headset in question) is a very quiet person, he did not talk much during the EVA, that is why we noticed the problem only when he tried to communicate the rover numbers to Hab at the end of the EVA when we were about to enter the Hab.
Solution: The problem does not look like it could be resolved apart from replacing the cable or the headset entirely. When radio number 5 has to go out on EVA, we will give him another headset to use.
Summary of GreenHab operations:
WATER USE: 13.4 gallons
Heater: Off
Supplemental light: Off
Harvest: cherry tomatoes 217 g, Basilic 4g, Mint 3g, Sage 0.5g
Summary of ScienceDome operations: extremophiles were put in cultures and incubated overnight. Arduino was programmed for pressure, altitude, temperature and humidity for a weather balloon.
Dual split: Heat or AC, Off
Summary of RAM operations: Assembly of 3D printer. Knobs on the oven were complicated to use so we worked a little bit on them and they work better now.
Summary of any observatory issues: see above
Summary of health and safety issues: Hole in the tunnel in front of the Science Dome because of wind gusts. Commander was blocked in there during repairs. Repairs were made by Sergii, everything nominal now.
Questions, concerns and requests to Mission Support:
Concerning water use, the water tank was filled today to 520 gallons during the day, and I measured the water level at the end of the day but not before the fill.

Copyright © The Mars Society. All rights reserved. | Main Site